A diffraction grating has 6 700 lines/cm. What is the slit separation? in M

a. 6.7E–3 m
b. 1.5E–6 m
c. 1.5E–9 m
d. 6.7E–7 m

I just want to make sure i understand the concept. You divide- 1/d and then convert m? Double-checking my work because im paranoid. Please write clearly :)
Answer #1

Basically, diffraction grating is a kind of instrument used to diffract the light into various components in different directions.

To find out the spacing between slits or slit seperation we just have to take inverse of number of lines per cm.

That is, slit separation d = 1 / number of lines per cm

Hence d = 1 / ( 6700 ) = 1.49 × 10-4 cm

But we need the separation in m

1 cm = 10-2 m

1.49 × 10-4 cm = 1.49 × 10-4 × 10-2 m

Therefore d = 1.49 × 10-6 m ( Answer)

Option B is correct answer.

Your concept is right.